OpenStudy (anonymous):

I need Algebra 1 help! The question is below. It was to fraction-y.

OpenStudy (anonymous):

What is the product of: \[\frac{ m^2 + 2m - 8 }{ 6p } \times \frac{ 12p }{ 4m + m^2 }\]

OpenStudy (anonymous):

okay let me see....

OpenStudy (anonymous):

Am I supposed to Cross Multiply?

OpenStudy (skullpatrol):

No.

OpenStudy (anonymous):

Good thing I didn't start cross multiplying.

OpenStudy (anonymous):

let me try to figure this out.I remember doing these problems before.

OpenStudy (anonymous):

okay,

OpenStudy (anonymous):

okay im still trying to figure it out.......i didn't forgot about ur question..:D

OpenStudy (anonymous):

multiply together the top multiply together the bottom

OpenStudy (anonymous):

if top and bottom has a similar multiple, then cancel it out

OpenStudy (anonymous):

example 12p= 2*6p top has a 6p bottom has a 6p, you can cancel it out because 6p/6p = 1

OpenStudy (anonymous):

uhh... that kinda confused me

OpenStudy (anonymous):

factor the numerator of the first term (m+4)(m-2)

OpenStudy (anonymous):

factor the denominator of the second term. (m)(m+4)

OpenStudy (anonymous):

Do you want the multiple choice answers?

OpenStudy (anonymous):

that would be better

OpenStudy (anonymous):

okay.. this might take a while. one second.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[\frac{ 16 }{ x-9 }, x \neq 1, 8 ,9 \] \[\frac{ 16 }{ x+8 }, x \neq -8, 1, 8\] \[\frac{ 16x - 65 }{ x^2 - 9x + 8 }, x \neq -8, -1\] \[\frac{ 16x - 65 }{ x^2 - 9x + 8 }, x \neq 1, 8 \]

OpenStudy (anonymous):

I just guessing, and I know it's wrong to guess. But, I think it's like D or something.

OpenStudy (anonymous):

well i don't know that exact answer but its okay to guess everyone does it

OpenStudy (anonymous):

Yeah, buts it's worth 10 points..

OpenStudy (anonymous):

okay...lets see...is this schoolwork??or homework??

OpenStudy (anonymous):

it's for Florida Virtual School, and I'm doing a review test and I'm having trouble with some of it

OpenStudy (anonymous):

okay let me see...

OpenStudy (anonymous):

look at the replys.There's an answer there.I don't know if its the right one.Just pick the one that is closes to your answer.okay?

OpenStudy (anonymous):

Urgh, I'm just gonna choose D.

OpenStudy (anonymous):

Okay.Good luck with your test!!Hop you pass!:)

OpenStudy (anonymous):

Yeah, if I don't I'll probably just ask my aunt for help.

OpenStudy (anonymous):

Wait, can I ask another question?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

it's not the right answer

OpenStudy (anonymous):

did u have another question vetdoc??

OpenStudy (anonymous):

the answer is not one of the options....

OpenStudy (anonymous):

yea i know...........idk what he is gonna do....

OpenStudy (anonymous):

i think thet are right, the answer makes sense

OpenStudy (anonymous):

he didnt wright the restrictions though

OpenStudy (anonymous):

can you halp me out on my post ?

OpenStudy (anonymous):